OpenStudy (anonymous):

How do I solve: 30 - 6y = 8

OpenStudy (nincompoop):

subtract 30 on both sides to isolate -6y - 30 + 30 - 6y = 8 - 30 = -6y = -22 look for y by dividing both sides by -6 -6y/-6 = -22/-6 = y = 22/6

sam (.sam.):

Its better to simplify 22/6

OpenStudy (nincompoop):

\[\frac{ 22 }{ 6 }= 3 \frac{ 2 }{ 3 }\]